A 3.0 amp current flowing through a solenoid, length 2.5 cm, produces a magnetic flux density in the air-filled core of 6.0 mT. How many complete turns of wire make up the solenoid?
Answer #1

magnetic field inside the solenoid is

B =mu*n*I

=mu*(N/L)*I

B*L=mu*N*I

N =(B*L)/(mu*I)

= (6*10^-3 T*0.025 m)/(4pi*10^-7*3 A)

= 39.7887 turns

= 39 turns